OpenStudy (anonymous):

Solve sqrt3 tanx - 1 = 0

OpenStudy (anonymous):

Okkk this one is pretty basic

OpenStudy (anonymous):

Move over the one to the other side and then divide by square root 3

OpenStudy (anonymous):

\[\sqrt[3]{tanx}-1=0 \]

OpenStudy (anonymous):

that's right ?

OpenStudy (anonymous):

nope its \[\sqrt{3} \tan x -1=0\]